Submit Quiz Quiz: Week 7 Quiz This Quiz: 17 pts 15 O Type here to search
Answer #1

So, triangles are similar

Formula:

AB/BD = AC/DC

x/9 = 5/15

x = 3

Also,

y/12 = 5/15

y = 4